OpenStudy (anonymous):

A solution of H2SO4(aq) with a molal concentration of 3.97 m has a density of 1.216 g/mL. What is the molar concentration of this solution?

OpenStudy (anonymous):

First Find the Mass of the Solute: 3.97 * 98 = 386.06 Nw Find the Total Volume of Solution 386.06 + 1216 = 1605.06 Nw Molarity = No of Moles of Solute / Volume ( 1 L) = 3 / 1.605 = 2.47 M